OpenStudy (anonymous):

If θ is between 0° and 90° and tanθ = 8/7, find cosθ A. 8√(113) B. 7√(113)/113 C. √(113)/7 D. 8√(113)/113

OpenStudy (anonymous):

SO i find it helpful to draw it out

OpenStudy (anonymous):



OpenStudy (anonymous):

Thats supposed to be theta Do you understand why the two legs are 8 and 7?

OpenStudy (anonymous):

Because that is tantheta

OpenStudy (anonymous):

Yes so you how can you find the hypotenuse of the triangle?

OpenStudy (anonymous):

Opposite adjacent something

OpenStudy (anonymous):

Umm try something simpler like pythagorean theorem to find the hypotnuse

OpenStudy (anonymous):

so do 8^2+7^2=c^2

OpenStudy (anonymous):

so c^2 = 15^2

OpenStudy (anonymous):

no you can't do that because you have to do the exponents first So do (8^2) THEN add (7^2)

OpenStudy (anonymous):

c^2 = 113?

OpenStudy (anonymous):

Yes now to get only c, you would take the square root and it would be sqrt(113)

OpenStudy (anonymous):

So now you would have 

OpenStudy (anonymous):

AS the third side

OpenStudy (anonymous):

so what is the formula for cos?

OpenStudy (anonymous):

The formula for cos is c2 = a2 + b2?

OpenStudy (anonymous):

Umm no, cosine is adjacent/hypotenuse right? SohCahToa, ring a bell?

OpenStudy (anonymous):

yes yes correct

OpenStudy (anonymous):

Right so now what is cosine of theta?

OpenStudy (anonymous):

According to the picture

OpenStudy (anonymous):

7/ √(113) ?

OpenStudy (anonymous):

Yup there you go :)

OpenStudy (anonymous):

So I guess my answer is B. then?

OpenStudy (anonymous):

Yes because that is what you get when you rationalize it

OpenStudy (anonymous):

Thank you good sir!

OpenStudy (anonymous):

No problem